# Video Event Reasoning and Prediction by Fusing World Knowledge from LLMs with Vision Foundation Models

### v1

@@ 3.1.2 Object-centric Feature Enhancement
While global features capture the overall scene dynamics, high-level reasoning often hinges on specific objects and their states. To provide the model with this structured information, we enhance our visual representation with object-centric features. We process keyframes from the video using a pre-trained, promptable segmentation model, the Segment Anything Model (SAM) (31). For each keyframe, we use an automated prompting mechanism (e.g., a grid-based point prompt) to generate a set of object masks $\{M_{j}\}_{j=1}^{K}$, where $K$ is the number of detected objects. For each object mask $M_{j}$, we use mask-based average pooling to extract its feature representation from the patch embeddings of the visual backbone. This results in a set of object tokens, $Z_{obj}$, for the entire video.
@@ 3.2 Vision-Language Fusion Core
A fundamental challenge in multimodal learning is bridging the "modality gap" between the continuous, high-dimensional space of visual features and the discrete, symbolic space of language. Simply projecting visual features into the language embedding space can be inefficient and may introduce noise. To address this, we employ a sophisticated fusion module inspired by the Q-Former architecture from BLIP-2 (33), which acts as an information bottleneck to distill the most relevant visual information for the LLM.
@@ 3.4.2 Stage 2: Instruction-based Fine-Tuning
After the fusion module is aligned with the two backbones, we move to the second stage to teach the model the higher-level reasoning and prediction tasks. In this stage, we use a curated, high-quality dataset of instruction-response pairs specific to video reasoning and prediction. We unfreeze the LLM parameters (or employ a parameter-efficient fine-tuning technique like LoRA (21)) and continue to train the Fusion Core. The objective remains a language modeling loss, but this time it is computed on the ground-truth reasoning or prediction text. This two-stage process ensures that the model first learns basic visual description before mastering complex cognitive tasks.
@@ 4.1.1 Training and Evaluation Datasets
Moving to the critical Stage 2 instruction fine-tuning, we address the scarcity of high-quality reasoning data by constructing a rich, hybrid dataset. We incorporate established academic benchmarks, including NExT-QA for its focus on causal and temporal reasoning, and ActivityNet-QA for its large scale and temporal query diversity. Recognizing the limitations of existing resources, the cornerstone of our fine-tuning data is our synthetically generated Causal-Vid-Instruct dataset. To create it, we prompted a powerful teacher model (GPT-4V) with 100,000 video clips sampled from diverse sources like Ego4D (11), generating detailed causal explanations and plausible future predictions. This synthetic data provides high-quality, targeted examples of the cognitive behaviors we aim to instill in our model.
Finally, to rigorously assess the generalization capabilities of our framework, we reserve several challenging benchmarks exclusively for zero-shot evaluation. These include the Test of Time (ToT) benchmark (39) to specifically diagnose temporal reasoning, the synthetic CLEVRER dataset (57) to probe for emergent physical and causal intuition, and the VCR benchmark (60) to evaluate commonsense reasoning on still frames derived from videos. None of these evaluation datasets were seen by the model during any phase of training.
@@ 4.1.2 Stage 2: Instruction Fine-Tuning Datasets
  • Causal-Vid-Instruct (Synthetic): High-quality, instruction-based video reasoning data is scarce. To overcome this, we generate a synthetic dataset. We sample 100,000 video clips from diverse datasets like Ego4D (11) and Something-Something v2. For each clip, we use a powerful proprietary teacher model (GPT-4V) to generate a "chain-of-thought" style causal explanation and a plausible future event prediction. The prompt given to the teacher model is: "Observe the following video clip. First, provide a step-by-step causal explanation of the events. Second, predict what is most likely to happen immediately after the clip ends." The resulting high-quality (video, instruction, response) triplets form our Causal-Vid-Instruct dataset, which is crucial for teaching the desired cognitive behaviors.
@@ 4.1.3 Evaluation-Only Benchmarks
  • Test of Time (ToT) (39): A recently proposed benchmark specifically designed to diagnose a model’s temporal reasoning abilities, including event ordering, duration comparison, and temporal localization. Its focus on challenging temporal queries makes it an ideal testbed for our reasoning claims.
  • CLEVRER (57): While a synthetic dataset, CLEVRER is a powerful diagnostic tool for testing causal and physical reasoning. It features videos of colliding objects where questions require understanding concepts like causality, object permanence, and collision dynamics. Success on this benchmark in a zero-shot setting would provide strong evidence of the model’s emergent physical intuition.
  • VCR (Visual Commonsense Reasoning) (60): We use the VCR benchmark to evaluate commonsense reasoning. For each video, we use the center frame and task the model with answering a challenging question and providing a rationale, testing its ability to transfer knowledge to an image-based reasoning context.
@@ 4.2.1 Model Architecture
Our framework is built upon powerful, publicly available foundation models. The Visual Perception Backbone is the InternVideo-B/16 model (52), which has demonstrated state-of-the-art performance on a wide range of video understanding tasks. The LLM-based Cognitive Reasoner is the instruction-tuned Llama-3-8B-Instruct model, known for its strong reasoning and language generation capabilities. Our Vision-Language Fusion Core is inspired by the Q-Former architecture and consists of 32 learnable queries, a hidden dimension of 768, and 8 cross-attention layers for distilling visual information.
@@ 4.2.2 Training Details
  • Stage 2 (Instruction Fine-Tuning): We fine-tune the model on our combined instruction dataset for 3 epochs. In this stage, we unfreeze the LLM and employ Parameter-Efficient Fine-Tuning using LoRA (21) to preserve its pre-trained knowledge while adapting it to our tasks. We set the LoRA rank $r=64$ and alpha $\alpha=128$, applying it to all linear layers of the LLM. The Fusion Core remains trainable. The learning rate is reduced to 2e-5, and we use a smaller batch size of 256. This stage was performed on 8 NVIDIA H100 GPUs for approximately 48 hours.
@@ 4.2.3 Baselines for Comparison
  • General Video-LLMs: We compare against leading video dialogue models, including Video-LLaMA (64) and Video-ChatGPT (38).
  • Reasoning-focused Models: For reasoning tasks, we compare against models designed for structured inference, such as SeViLA (59) and the tool-augmented ViperGPT (45).
  • Prediction-focused Models: For future prediction, we compare against a strong baseline for action anticipation from (1) and adapt a video generation model, MCVD (49), to generate textual descriptions of its predicted future frames.
@@ 5.1 Main Quantitative Comparison
We begin by assessing the overall performance of our model against existing methods on established benchmarks for video reasoning and open-ended generation. The results, summarized in Table LABEL:tab:reasoning_results and Table LABEL:tab:prediction_results, consistently demonstrate the superiority of our approach.
@@ 5.1.1 Performance on Video Reasoning Tasks
As shown in Table LABEL:tab:reasoning_results, on reasoning-centric, multiple-choice QA benchmarks like NExT-QA and VCR, our model sets a new state of the art. On NExT-QA, which evaluates causal and temporal understanding, our model achieves a top accuracy score, significantly outperforming general video-LLMs like Video-LLaMA (64) and Video-ChatGPT (38). We attribute this performance gain to two key factors. First, our explicit inclusion of object-centric features provides the model with a more structured and grounded representation of entity interactions, which is often crucial for answering "why" questions. Second, our use of the synthetically generated Causal-Vid-Instruct dataset during Stage 2 fine-tuning directly exposes the model to the patterns of causal language, a significant advantage over models trained primarily on descriptive captions.
Compared to reasoning-focused models like SeViLA (59), which employs a self-chained reasoning process, our model’s performance suggests that fusing a powerful, pre-trained LLM with rich visual inputs allows it to perform implicit reasoning more effectively than explicitly decomposing the problem. Furthermore, our framework outperforms ViperGPT (45), an innovative approach that uses an LLM to generate code. While ViperGPT is powerful for queries that can be answered by composing existing vision tools, our model excels at tasks requiring holistic, commonsense understanding of unscripted events, which cannot be easily solved by a sequence of API calls.
@@ 5.1.2 Performance on Open-ended Generation and Prediction
The true test of our model’s cognitive capabilities lies in its ability to generate free-form, coherent text for reasoning and prediction. In Table LABEL:tab:prediction_results, we report results on our open-ended test set using a suite of metrics. In terms of n-gram-based scores (BLEU, ROUGE-L, CIDEr), our model is highly competitive, indicating its fluency in generating grammatically correct and relevant text. However, these metrics are known to be limited. On semantic similarity metrics like BERTScore, our model shows a more pronounced advantage, confirming that its generations are not just syntactically similar but also semantically closer to the ground truth.
@@ 5.2 Zero-Shot Generalization Performance
A key objective of our work is to build a model that learns generalizable reasoning skills rather than simply overfitting to the patterns in the fine-tuning data. To assess this, we evaluated our model on two challenging benchmarks, CLEVRER and Test of Time (ToT), without any task-specific training. The results are presented in Table LABEL:tab:zeroshot_results.
On the CLEVRER dataset (57), which tests physical and causal reasoning in a synthetic 3D environment, our model achieves a surprisingly high accuracy in a zero-shot setting. It significantly outperforms all video-dialogue baselines, which often fail to grasp the underlying physics of collisions and object permanence. This suggests that the combination of large-scale video pre-training and the LLM’s inherent (though imperfect) knowledge of physics allows our model to develop an emergent intuition for physical laws.
Similarly, on the Test of Time (ToT) benchmark (39), which is designed to diagnose temporal reasoning, our model demonstrates a strong zero-shot capability. It successfully answers complex questions about event ordering, duration, and relationships, far exceeding the performance of models not explicitly designed for such fine-grained temporal analysis. This success indicates that our model has learned abstract principles of time and sequence from its training, rather than relying on dataset-specific cues. This ability to generalize is critical for real-world applications where systems must constantly encounter and interpret novel scenarios.
@@ 5.3 Ablation Studies
To dissect our framework and quantitatively validate our architectural choices, we conducted a series of in-depth ablation studies, with the results summarized in Table LABEL:tab:ablation_studies. The findings unequivocally confirm that each component plays a synergistic and indispensable role. The most critical component is our Vision-Language Fusion Core; replacing our Q-Former-inspired module with a simpler mean-pooling and linear projection approach (‘w/o Fusion Core‘) led to a catastrophic degradation in performance. This demonstrates that the fusion module is not merely a projector but a vital information bottleneck that effectively filters and translates visual data for the LLM. Furthermore, the importance of a structured visual input was confirmed when we removed the object-centric features (‘w/o Object Features‘). This resulted in a notable drop in performance on tasks requiring fine-grained causal reasoning about entity interactions, validating our hybrid feature extraction strategy. The necessity of our two-stage training protocol was also made evident. The model trained only on alignment (‘w/o Stage 2 FT‘) could produce basic descriptions but failed entirely at reasoning tasks, highlighting that instruction fine-tuning is essential for eliciting cognitive behaviors. Similarly, training without our synthetic Causal-Vid-Instruct dataset (‘w/o Synthetic Data‘) yielded a model with weaker explanatory and predictive power, confirming the value of high-quality, targeted instruction data. Finally, replacing our LLM with a smaller backbone (‘Smaller LLM‘) resulted in less coherent and nuanced reasoning, reinforcing the conclusion that the framework’s ultimate capability is tightly coupled with the power of its cognitive core.
5.4 Qualitative Analysis and Visualization
Quantitative metrics provide a measure of overall performance, but qualitative examples offer deeper insights into the model’s strengths and weaknesses. We present several examples in Fig. 2 and Fig. 2.
@@ 5.4 Qualitative Analysis and Visualization
In a successful example of causal reasoning, shown in Fig. LABEL:fig:qual_success(a), the input video depicts a person accidentally knocking over a plant pot while reaching for a book on a high shelf. A baseline model provides a simple description: "A person reaches for a book and a pot falls." In contrast, our model generates a detailed, causal explanation: "The person stretches to reach the top shelf, causing their elbow to brush against the plant pot. The contact provides enough force to unbalance the pot, which then tips over the edge and falls to the ground due to gravity." This demonstrates our model’s ability to construct a complete causal chain from observed events.
In another example focused on prediction (Fig. LABEL:fig:qual_success(b)), the video shows a chef meticulously seasoning a piece of steak and placing it in a hot, oiled pan. Our model predicts: "The chef will likely sear the steak for a couple of minutes on each side to create a crust before potentially placing it in an oven to finish cooking." This prediction is not only accurate but also demonstrates domain-specific knowledge (about searing and finishing steaks) that is clearly drawn from the LLM’s world knowledge.
@@
Prompt: “Given what has occurred in this video, what is the most likely subsequent event?”
Baseline Model Output:
The person will leave the house.
Our Model Output:
The person is most likely about to go for a run or workout …
Figure 4: (b) Example of future prediction. Our model …
However, our model is not infallible. Fig. 2 illustrates a common failure mode: factual hallucination. In a video showing a mechanic changing a tire, our model correctly identifies most steps but hallucinates a detail: "…after tightening the lug nuts, the mechanic uses a torque wrench to ensure they are at the specified tightness." While this is the correct professional procedure, a torque wrench was not visible or used in the video clip. This type of error, as investigated in works like (58), occurs when the LLM’s powerful prior knowledge overrides the immediate visual evidence. It highlights the ongoing challenge of achieving perfect and robust visual grounding, which remains a key area for future research.
@@ 5.5 Discussion and Limitations
The collective results present strong evidence that our framework represents a significant step forward in video understanding. […] Despite these promising results, we acknowledge several limitations that chart clear directions for future research. A primary consideration is the substantial computational cost associated with training and inference, a common challenge for foundation models. Future work should explore advanced model compression techniques (7) and more efficient architectures to democratize access to these capabilities. Moreover, our model’s advanced reasoning is partly derived from our synthetically generated instruction dataset, creating a dependency on the teacher model and a potential for inheriting its latent biases. Developing methods for fine-tuning with more diverse, human-curated data or through direct reinforcement learning from human feedback would be a valuable next step. The persistent issue of factual grounding and hallucination, while mitigated, also warrants further attention. As our qualitative analysis shows, ensuring every statement is perfectly grounded in visual evidence remains an open problem, suggesting a need for tighter fusion mechanisms or post-hoc verification modules. Finally, a notable gap remains between current benchmarks and the unstructured, long-form complexity of real-world video. Scaling our methods to handle hour-long inputs and reason about ambiguous social dynamics, such as those found in datasets like Ego-Exo4D (11), will be a major and exciting challenge for the field.

### v2

@@ 3.1.2 Object-centric Feature Enhancement
While global features capture the overall scene dynamics, high-level reasoning often hinges on specific objects and their states. To provide the model with this structured information, we enhance our visual representation with object-centric features. We process keyframes from the video using a pre-trained, promptable segmentation model, the Segment Anything Model (SAM) (kirillov2023segment). For each keyframe, we use an automated prompting mechanism (e.g., a grid-based point prompt) to generate a set of object masks $\{M_{j}\}_{j=1}^{K}$, where $K$ is the number of detected objects. For each object mask $M_{j}$, we use mask-based average pooling to extract its feature representation from the patch embeddings of the visual backbone. This results in a set of object tokens, $Z_{obj}$, for the entire video.
@@ 3.2 Vision-Language Fusion Core
A fundamental challenge in multimodal learning is bridging the "modality gap" between the continuous, high-dimensional space of visual features and the discrete, symbolic space of language. Simply projecting visual features into the language embedding space can be inefficient and may introduce noise. To address this, we employ a sophisticated fusion module inspired by the Q-Former architecture from BLIP-2 (li2023blip), which acts as an information bottleneck to distill the most relevant visual information for the LLM.
@@ 3.4.2 Stage 2: Instruction-based Fine-Tuning
After the fusion module is aligned with the two backbones, we move to the second stage to teach the model the higher-level reasoning and prediction tasks. In this stage, we use a curated, high-quality dataset of instruction-response pairs specific to video reasoning and prediction. We unfreeze the LLM parameters (or employ a parameter-efficient fine-tuning technique like LoRA (hu2021lora)) and continue to train the Fusion Core. The objective remains a language modeling loss, but this time it is computed on the ground-truth reasoning or prediction text. This two-stage process ensures that the model first learns basic visual description before mastering complex cognitive tasks.
@@ 4.1.1 Training and Evaluation Datasets
Moving to the critical Stage 2 instruction fine-tuning, we address the scarcity of high-quality reasoning data by constructing a rich, hybrid dataset. We incorporate established academic benchmarks, including NExT-QA for its focus on causal and temporal reasoning, and ActivityNet-QA for its large scale and temporal query diversity. Recognizing the limitations of existing resources, the cornerstone of our fine-tuning data is our synthetically generated Causal-Vid-Instruct dataset. To create it, we prompted a powerful teacher model (GPT-4V) with 100,000 video clips sampled from diverse sources like Ego4D (gandhi2023egoexo4d), generating detailed causal explanations and plausible future predictions. This synthetic data provides high-quality, targeted examples of the cognitive behaviors we aim to instill in our model.
Finally, to rigorously assess the generalization capabilities of our framework, we reserve several challenging benchmarks exclusively for zero-shot evaluation. These include the Test of Time (ToT) benchmark (momeni2023test) to specifically diagnose temporal reasoning, the synthetic CLEVRER dataset (yi2020clevrer) to probe for emergent physical and causal intuition, and the VCR benchmark (zellers2019vcr) to evaluate commonsense reasoning on still frames derived from videos. None of these evaluation datasets were seen by the model during any phase of training.
@@ 4.1.2 Stage 2: Instruction Fine-Tuning Datasets
  • Causal-Vid-Instruct (Synthetic): High-quality, instruction-based video reasoning data is scarce. To overcome this, we generate a synthetic dataset. We sample 100,000 video clips from diverse datasets like Ego4D (gandhi2023egoexo4d) and Something-Something v2. For each clip, we use a powerful proprietary teacher model (GPT-4V) to generate a "chain-of-thought" style causal explanation and a plausible future event prediction. The prompt given to the teacher model is: "Observe the following video clip. First, provide a step-by-step causal explanation of the events. Second, predict what is most likely to happen immediately after the clip ends." The resulting high-quality (video, instruction, response) triplets form our Causal-Vid-Instruct dataset, which is crucial for teaching the desired cognitive behaviors.
@@ 4.1.3 Evaluation-Only Benchmarks
  • Test of Time (ToT) (momeni2023test): A recently proposed benchmark specifically designed to diagnose a model’s temporal reasoning abilities, including event ordering, duration comparison, and temporal localization. Its focus on challenging temporal queries makes it an ideal testbed for our reasoning claims.
  • CLEVRER (yi2020clevrer): While a synthetic dataset, CLEVRER is a powerful diagnostic tool for testing causal and physical reasoning. It features videos of colliding objects where questions require understanding concepts like causality, object permanence, and collision dynamics. Success on this benchmark in a zero-shot setting would provide strong evidence of the model’s emergent physical intuition.
  • VCR (Visual Commonsense Reasoning) (zellers2019vcr): We use the VCR benchmark to evaluate commonsense reasoning. For each video, we use the center frame and task the model with answering a challenging question and providing a rationale, testing its ability to transfer knowledge to an image-based reasoning context.
@@ 4.2.1 Model Architecture
Our framework is built upon powerful, publicly available foundation models. The Visual Perception Backbone is the InternVideo-B/16 model (wang2022internvideo), which has demonstrated state-of-the-art performance on a wide range of video understanding tasks. The LLM-based Cognitive Reasoner is the instruction-tuned Llama-3-8B-Instruct model, known for its strong reasoning and language generation capabilities. Our Vision-Language Fusion Core is inspired by the Q-Former architecture and consists of 32 learnable queries, a hidden dimension of 768, and 8 cross-attention layers for distilling visual information.
@@ 4.2.2 Training Details
  • Stage 2 (Instruction Fine-Tuning): We fine-tune the model on our combined instruction dataset for 3 epochs. In this stage, we unfreeze the LLM and employ Parameter-Efficient Fine-Tuning using LoRA (hu2021lora) to preserve its pre-trained knowledge while adapting it to our tasks. We set the LoRA rank $r=64$ and alpha $\alpha=128$, applying it to all linear layers of the LLM. The Fusion Core remains trainable. The learning rate is reduced to 2e-5, and we use a smaller batch size of 256. This stage was performed on 8 NVIDIA H100 GPUs for approximately 48 hours.
@@ 4.2.3 Baselines for Comparison
  • General Video-LLMs: We compare against leading video dialogue models, including Video-LLaMA (zhang2023videollama) and Video-ChatGPT (maaz2023videochatgpt).
  • Reasoning-focused Models: For reasoning tasks, we compare against models designed for structured inference, such as SeViLA (yu2023sevila) and the tool-augmented ViperGPT (suris2023viperGPT).
  • Prediction-focused Models: For future prediction, we compare against a strong baseline for action anticipation from (abu-farha-2021-time) and adapt a video generation model, MCVD (voleti2022mcvd), to generate textual descriptions of its predicted future frames.
@@ 5.1.1 Performance on Video Reasoning Tasks
On reasoning-centric, multiple-choice QA benchmarks like NExT-QA and VCR, our model sets a new state of the art. On NExT-QA, which evaluates causal and temporal understanding, our model achieves a top accuracy score, significantly outperforming general video-LLMs like Video-LLaMA (zhang2023videollama) and Video-ChatGPT (maaz2023videochatgpt). We attribute this performance gain to two key factors. First, our explicit inclusion of object-centric features provides the model with a more structured and grounded representation of entity interactions, which is often crucial for answering "why" questions. Second, our use of the synthetically generated Causal-Vid-Instruct dataset during Stage 2 fine-tuning directly exposes the model to the patterns of causal language, a significant advantage over models trained primarily on descriptive captions.
Compared to reasoning-focused models like SeViLA (yu2023sevila), which employs a self-chained reasoning process, our model’s performance suggests that fusing a powerful, pre-trained LLM with rich visual inputs allows it to perform implicit reasoning more effectively than explicitly decomposing the problem. Furthermore, our framework outperforms ViperGPT (suris2023viperGPT), an innovative approach that uses an LLM to generate code. While ViperGPT is powerful for queries that can be answered by composing existing vision tools, our model excels at tasks requiring holistic, commonsense understanding of unscripted events, which cannot be easily solved by a sequence of API calls.
@@ 5.1.2 Performance on Open-ended Generation and Prediction
The true test of our model’s cognitive capabilities lies in its ability to generate free-form, coherent text for reasoning and prediction. we report results on our open-ended test set using a suite of metrics. In terms of n-gram-based scores (BLEU, ROUGE-L, CIDEr), our model is highly competitive, indicating its fluency in generating grammatically correct and relevant text. However, these metrics are known to be limited. On semantic similarity metrics like BERTScore, our model shows a more pronounced advantage, confirming that its generations are not just syntactically similar but also semantically closer to the ground truth.
@@ 5.2 Zero-Shot Generalization Performance
A key objective of our work is to build a model that learns generalizable reasoning skills rather than simply overfitting to the patterns in the fine-tuning data. To assess this, we evaluated our model on two challenging benchmarks, CLEVRER and Test of Time (ToT), without any task-specific training.
On the CLEVRER dataset (yi2020clevrer), which tests physical and causal reasoning in a synthetic 3D environment, our model achieves a surprisingly high accuracy in a zero-shot setting. It significantly outperforms all video-dialogue baselines, which often fail to grasp the underlying physics of collisions and object permanence. This suggests that the combination of large-scale video pre-training and the LLM’s inherent (though imperfect) knowledge of physics allows our model to develop an emergent intuition for physical laws.
Similarly, on the Test of Time (ToT) benchmark (momeni2023test), which is designed to diagnose temporal reasoning, our model demonstrates a strong zero-shot capability. It successfully answers complex questions about event ordering, duration, and relationships, far exceeding the performance of models not explicitly designed for such fine-grained temporal analysis. This success indicates that our model has learned abstract principles of time and sequence from its training, rather than relying on dataset-specific cues. This ability to generalize is critical for real-world applications where systems must constantly encounter and interpret novel scenarios.
@@ 5.3 Ablation Studies
The findings unequivocally confirm that each component plays a synergistic and indispensable role. The most critical component is our Vision-Language Fusion Core; replacing our Q-Former-inspired module with a simpler mean-pooling and linear projection approach (‘w/o Fusion Core‘) led to a catastrophic degradation in performance. This demonstrates that the fusion module is not merely a projector but a vital information bottleneck that effectively filters and translates visual data for the LLM. Furthermore, the importance of a structured visual input was confirmed when we removed the object-centric features (‘w/o Object Features‘). This resulted in a notable drop in performance on tasks requiring fine-grained causal reasoning about entity interactions, validating our hybrid feature extraction strategy. The necessity of our two-stage training protocol was also made evident. The model trained only on alignment (‘w/o Stage 2 FT‘) could produce basic descriptions but failed entirely at reasoning tasks, highlighting that instruction fine-tuning is essential for eliciting cognitive behaviors. Similarly, training without our synthetic Causal-Vid-Instruct dataset (‘w/o Synthetic Data‘) yielded a model with weaker explanatory and predictive power, confirming the value of high-quality, targeted instruction data. Finally, replacing our LLM with a smaller backbone (‘Smaller LLM‘) resulted in less coherent and nuanced reasoning, reinforcing the conclusion that the framework’s ultimate capability is tightly coupled with the power of its cognitive core.
@@
However, our model is not infallible. Fig.  illustrates a common failure mode: factual hallucination. In a video showing a mechanic changing a tire, our model correctly identifies most steps but hallucinates a detail: "…after tightening the lug nuts, the mechanic uses a torque wrench to ensure they are at the specified tightness." While this is the correct professional procedure, a torque wrench was not visible or used in the video clip. This type of error, as investigated in works like (yin2023woodpecker), occurs when the LLM’s powerful prior knowledge overrides the immediate visual evidence. It highlights the ongoing challenge of achieving perfect and robust visual grounding, which remains a key area for future research.
@@ 5.4 Discussion and Limitations
The collective results present strong evidence that our framework represents a significant step forward in video understanding. […] Despite these promising results, we acknowledge several limitations that chart clear directions for future research. A primary consideration is the substantial computational cost associated with training and inference, a common challenge for foundation models. Future work should explore advanced model compression techniques and more efficient architectures to democratize access to these capabilities. Moreover, our model’s advanced reasoning is partly derived from our synthetically generated instruction dataset, creating a dependency on the teacher model and a potential for inheriting its latent biases. Developing methods for fine-tuning with more diverse, human-curated data or through direct reinforcement learning from human feedback would be a valuable next step. The persistent issue of factual grounding and hallucination, while mitigated, also warrants further attention. As our qualitative analysis shows, ensuring every statement is perfectly grounded in visual evidence remains an open problem, suggesting a need for tighter fusion mechanisms or post-hoc verification modules. Finally, a notable gap remains between current benchmarks and the unstructured, long-form complexity of real-world video. Scaling our methods to handle hour-long inputs and reason about ambiguous social dynamics, such as those found in datasets like Ego-Exo4D (gandhi2023egoexo4d), will be a major and exciting challenge for the field.